27) If a company has a product with a margin of 74% and the plant is running at 90% capacity, a discount of 25% on marginal s thanks
Answer #1

Margin = 74%

Plant capacity used = 90%

Discount on marginal sales = 25%

Since the margin is more than discount on marginal sales, hence a discount of 25% on marginal sales will still result in profit and thus it is acceptable.

It can be understood as under:

Let the sales at 100% Capacity utilization of plant be = $100,000

Sales at 90% capacity = 100,000 x 90%

= $90,000

Sales for unused capacity of plant = 100,000-90,000

= $10,000

Since margin is 74%, hence cost of goods sold is 26%

Cost of marginal sales = 10,000 x 26%

= $2,600

Discount on marginal sales = 25%

= 10,000 x 25%

= $2,500

Net Sales = Sales - Discount

= 10,000-2,500

= $7,500

Gross margin on marginal sales = Net Sales - Cost of marginal sales

= 7,500-2,600

= $4,900

Hence, even after 25% sales discount on marginal sales, it will result in profit. Hence, 25% discount on marginal sales is acceptable.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks